deathsindragon:

bailee had a gross income of $2358.33 during each pay period in 2009. if she got payed monthly, how much of her pay was deducted for FICA in 2009? A. $1881.98 B. $ 410.35 C. $1754.60 D. $2164.95

WendyMorales2007:

$2358.33 * 12 months = $28,299.96 then subtract 28,299.96- 2358.33 btw im not sure if im right

Hoodmemes:

U were some what on the right track but not correct

deathsindragon:

no, it gave me 25940.67 when i subtracted those 2

Hoodmemes:

Im thinking u would wanna multply her income in total by the months in a year which would look like \[2358.33 \times 12 = 28299.96.\] I aint sure what u would do after that so give me like 14 minutes to work that out

deathsindragon:

mk

Hoodmemes:

We need to do a equation wit them percentages

addisonwk:

Wendy is right

addisonwk:

no nvm

Hoodmemes:

So add the percentages together like \[6.2+1.45 = 7.65\] the divide that answer by 100 which equals \[7.65 \times 100=0.0765\]

Hoodmemes:

Then u would multiply which looks like dis \[28299.96\times 0.0765 = (blank)\] Do that and u get your answer

Hoodmemes:

9u understand what im sayin?)

deathsindragon:

ye

Hoodmemes:

so whats your final answer

deathsindragon:

2164.94

deathsindragon:

but if i round it its 2164.95

Hoodmemes:

correct

Hoodmemes:

@Hoodmemes

Hoodmemes:

Not me putting a multiplication sign instead of a division sign when I clearly said "divide"